# Lambda Expected Shortfall

### v1

@@ 2.3 Expected Shortfall
Hence, we have $\rho(X)\geqslant\mathbb{E}[X]$ in the same sense as the argument in (i). As $\mathbb{E}$ dominates the essential infimum $\mathrm{VaR}^{+}_{0}$, it implies that $\mathbb{E}$ is the smallest quasi-convex and law-invariant risk measure that dominates $\mathrm{VaR}^{+}_{0}$. The proof is complete. ∎
Theorem 1 is stronger than two classic results: Föllmer and Schied (2016, Theorem 4.67), which requires $\rho$ in (7) to be convex, monetary and Fatou-continuous, and Delbaen (2012, Theorem 52), which requires $\rho$ in (7) to be coherent. [^3] We say a risk measure $\rho:\mathcal{X}\to\mathbb{R}$ is Fatou-continuous if it is lower semicontinuous under bounded pointwise convergence: For all bounded $X,X_{1},X_{2},\dots\in\mathcal{X}$ such that $X_{n}\to X$ pointwise as $n\to\infty$, $\rho(X)\leqslant\liminf_{n\to\infty}\rho(X_{n})$. Both of the two results above further assumed that $\rho$ takes finite values and $\alpha\in(0,1)$, but these differences are not essential.
We note that (7) fails for $\alpha=0$ because $\mathrm{VaR}_{0}=$-$\infty$ is quasi-convex, and the smallest quasi-convex and law-invariant risk measure dominating $\mathrm{VaR}_{0}$ is itself instead of $\mathrm{ES}_{0}=\mathbb{E}$. Similarly, (8) fails for $\alpha=1$ because $\mathrm{VaR}_{1}^{+}=\infty$ is quasi-convex, and the smallest quasi-convex and law-invariant risk measure dominating $\mathrm{VaR}_{1}^{+}$ is itself instead of $\mathrm{ES}_{1}=\mathrm{VaR}_{1}$.
For $\alpha\in(0,1)$, the statement in Theorem 1 extends directly to $L^{1}$, but for $\alpha=0$ and $\alpha=1$, some minor adjustments are needed, which we discuss in Section 6.
@@ 3 Lambda ES
The main question of this paper is how to define a counterpart to $\Lambda$-VaR that is similar to ES as a counterpart to VaR. We will call this counterpart $\Lambda$-Expected Shortfall, or $\Lambda$-ES for short.
3.1 Requirements for Lambda ES
By defining a new class of risk measures, there should be some clear gain. Otherwise, the newly defined class is not useful. The following properties are also satisfied by $\Lambda$-VaR, and they will be considered as basic requirements for $\Lambda$-ES. We believe their desirability is self-evident.
@@ 3.1 Requirements for Lambda ES
The next four properties are additional crucial requirements for $\Lambda$-ES to be considered a useful alternative to $\Lambda$-VaR, and they highlight the features of ES over VaR.
@@
  • $\Lambda$-ES should be $L^{1}$-continuous. This guarantees the robustness of $\Lambda$-ES with respect to slight changes to random losses, sharing the same advantage as ES and other coherent risk measures in risk management practice.
Some other properties, such as normalization, cash subadditivity, and quasi-concavity in mixtures are also desirable and natural from the corresponding properties of ES, but they may be less crucial as a criterion. They will all be satisfied by our proposed candidate for $\Lambda$-ES.
3.2 A formal definition of Lambda ES
For a decreasing function $\Lambda:\mathbb{R}\to[0,1]$, to define $\Lambda$-ES using the idea of Theorem 1, the formulation is obvious:
$\min\{\rho:L^{\infty}\to\overline{\mathbb{R}}\mid\rho\geqslant\mathrm{VaR}_{\Lambda}~\mbox{and $\rho$ is quasi-convex and law-invariant}\}.$ (9)
Theorem 1 of Han et al. (2025) gives a representation result of $\Lambda$-VaR. Below, we restate the result and extend it to $\Lambda$-$\mathrm{VaR}^{+}$.
Proposition 1.
Let $\Lambda:\mathbb{R}\to[0,1]$ be a decreasing function. The risk measures $\mathrm{VaR}_{\Lambda}$ and $\mathrm{VaR}^{+}_{\Lambda}$ admit the following representations.
Proof.
Equation (10) holds directly by Theorem 1 of Han et al. (2025) for all decreasing functions $\Lambda:\mathbb{R}\to[0,1]$ that are not constantly 0. For $\Lambda\equiv 0$ (meaning that $\Lambda$ is constantly 0), we have
Thus (10) holds for all decreasing functions $\Lambda:\mathbb{R}\to[0,1]$.
The proof of (11) is in the same spirit as that in Theorem 1 of Han et al. (2025) to prove (10). We write it below for completeness. For a decreasing function $\Lambda:\mathbb{R}\to[0,1]$ and $X\in L^{0}$, we have
$=\sup\{x\in\mathbb{R}:\mathrm{VaR}^{+}_{\Lambda(x)}(X)\geqslant x\}$
$=\sup\{\mathrm{VaR}^{+}_{\Lambda(x)}(X)\wedge x:\mathrm{VaR}_{\Lambda(x)}(X)\geqslant x\}\leqslant\sup_{x\in\mathbb{R}}\left\{\mathrm{VaR}^{+}_{\Lambda(x)}(X)\wedge x\right\},$
and similarly,
$\mathrm{VaR}^{+}_{\Lambda}(X)$ $=\sup\{x\in\mathbb{R}:\mathrm{VaR}^{+}_{\Lambda(x)}(X)\geqslant x\}$
$=\inf\{x\in\mathbb{R}:\mathrm{VaR}^{+}_{\Lambda(x)}(X)<x\}$
$=\inf\{\mathrm{VaR}^{+}_{\Lambda(x)}(X)\vee x:\mathrm{VaR}_{\Lambda(x)}(X)<x\}\geqslant\inf_{x\in\mathbb{R}}\left\{\mathrm{VaR}^{+}_{\Lambda(x)}(X)\vee x\right\}.$
Therefore, we have
$\mathrm{VaR}^{+}_{\Lambda}(X)\leqslant\sup_{x\in\mathbb{R}}\left\{\mathrm{VaR}^{+}_{\Lambda(x)}(X)\wedge x\right\}\leqslant\inf_{x\in\mathbb{R}}\left\{\mathrm{VaR}^{+}_{\Lambda(x)}(X)\vee x\right\}\leqslant\mathrm{VaR}^{+}_{\Lambda}(X).$
The proof is complete. ∎
Inspired by (10), we formally define the Lambda Expected Shortfall as follows.
@@ 3.2 A formal definition of Lambda ES
Figure 1 illustrates the definition of $\Lambda$-ES when the function $\Lambda$ is continuous and discontinuous, respectively. Writing $x^{*}=\mathrm{ES}_{\Lambda}(X)$, we can see that $(x^{*},x^{*})$ is the unique intersection point between the graph (linearly interpolated) of the function $x\mapsto\mathrm{ES}_{\Lambda(x)}(X)$ and the graph of the identity. From the right panel of Figure 1, we can also see that whether $x\mapsto\mathrm{ES}_{\Lambda(x)}(X)$ is left- or right-continuous at $x^{*}$ (or neither) does not matter.
@@
By definition, $\mathrm{ES}_{\Lambda}$ is finite on $L^{\infty}$; see also Proposition 2 (a) below. The next result shows that $\mathrm{ES}_{\Lambda}$ in Definition 1 satisfies (9), generalizing the fact that $\mathrm{ES}_{\alpha}$ satisfies (7). Moreover, it admits an alternative representation (13), which can be clearly seen from Figure 1.
@@
    For a decreasing function $\Lambda:\mathbb{R}\to(0,1]$, the smallest quasi-convex and law-invariant risk measure on $L^{\infty}$ dominating $\mathrm{VaR}_{\Lambda}$ is $\mathrm{ES}_{\Lambda}$, that is, ES_Λ= min{ρ: L^∞→Rρ⩾VaR_Λ and $\rho$ is quasi-convex and law-invariant}.
@@
    For a decreasing function $\Lambda:\mathbb{R}\to[0,1)$, the smallest quasi-convex and law-invariant risk measure on $L^{\infty}$ dominating $\mathrm{VaR}^{+}_{\Lambda}$ is $\mathrm{ES}_{\Lambda}$, that is, ES_Λ= min{ρ: L^∞→Rρ⩾VaR^+_Λ and $\rho$ is quasi-convex and law-invariant}.
@@
Using (10)–(12), we can see that $\mathrm{ES}_{\Lambda}$ dominates $\mathrm{VaR}_{\Lambda}$ for $\Lambda:\mathbb{R}\to[0,1]$ (resp. $\mathrm{VaR}^{+}_{\Lambda}$ for $\Lambda:\mathbb{R}\to[0,1)$) because $\mathrm{ES}_{\alpha}\geqslant\mathrm{VaR}_{\alpha}$ for all $\alpha\in[0,1]$ (resp. $\mathrm{ES}_{\alpha}\geqslant\mathrm{VaR}^{+}_{\alpha}$ for all $\alpha\in[0,1)$). Moreover, $\mathrm{ES}_{\Lambda}$ is law-invariant by definition. Next, we show that $\mathrm{ES}_{\Lambda}$ is quasi-convex. Note that for any given $\alpha\in[0,1]$, $\mathrm{ES}_{\alpha}$ is quasi-convex. Further, an increasing transform of a quasi-convex function is quasi-convex, as well as the supremum of a set of quasi-convex functions. Using these facts,
$\mathrm{ES}_{\Lambda(x)}\mbox{ is quasi-convex for each $x\in\mathbb{R}$}$
$\implies\mathrm{ES}_{\Lambda(x)}\wedge x\mbox{ is quasi-convex for each $x\in\mathbb{R}$}$
@@
Next, for $\Lambda:\mathbb{R}\to(0,1]$, we show that for any $\rho$ that is quasi-convex, law-invariant, and satisfying $\rho\geqslant\mathrm{VaR}_{\Lambda}$, it must be $\rho\geqslant\mathrm{ES}_{\Lambda}$. For any $X\in L^{\infty}$, we have:
@@
$\implies\mbox{for all $x\in\mathbb{R}:~$}\rho(X)\geqslant\left(\mathrm{VaR}_{\Lambda(x)}(X)\wedge x\right)$
@@
For $\Lambda:\mathbb{R}\to[0,1)$, for any $\rho$ that is quasi-convex, law-invariant, and satisfying $\rho\geqslant\mathrm{VaR}^{+}_{\Lambda}$, we have $\rho(X)\geqslant\mathrm{ES}_{\Lambda}(X)$ for any $X\in L^{\infty}$ with the same argument as above by replacing $\mathrm{VaR}_{\Lambda(x)}$ by $\mathrm{VaR}^{+}_{\Lambda(x)}$ for all $x\in\mathbb{R}$. This completes the proof of statements (i) and (ii). The final statement in (13) follows from (10), by noting that an ES curve $\alpha\mapsto\mathrm{ES}_{\alpha}(X)$ for $X\in L^{\infty}$ can be written as a VaR (resp. $\mathrm{VaR}^{+}$) curve $\alpha\mapsto\mathrm{VaR}_{\alpha}(Y)$ for some $Y\in L^{0}$ on $\alpha\in(0,1]$ (resp. $\alpha\in[0,1)$); see e.g., Lemma 4.5 of Burzoni et al. (2022). ∎
In Theorem 2, the reason to exclude 0 in part (i) and $1$ in part (ii) from the range of $\Lambda$ is the same as that in Theorem 1, where 0 (resp. $1$) is excluded from the domination of $\mathrm{ES}_{\alpha}$ over $\mathrm{VaR}_{\alpha}$ (resp. $\mathrm{VaR}_{\alpha}^{+}$).
@@
moreover, if $\Lambda$ is continuous, then
$\mathrm{ES}_{\Lambda(x)}(X)=x~\iff~\mathrm{ES}_{\Lambda}(X)=x.$ (16)
The relations (14)–(16) will be convenient in some proof arguments.
@@
It is clear that $\mathrm{ES}_{\Lambda}$ is parameterized only by the function $\Lambda$, and $\mathrm{ES}_{\Lambda}=\mathrm{ES}_{\alpha}$ when $\Lambda\equiv\alpha$ for some $\alpha\in(0,1)$. It satisfies several other desirable properties as a good candidate for $\Lambda$-$\mathrm{ES}$ as discussed in Section 3.1, which we summarize in the following result.
@@
For any decreasing function $\Lambda:\mathbb{R}\to[0,1]$, the risk measure $\mathrm{ES}_{\Lambda}$ satisfies the following properties: (a) $\mathrm{ES}_{\Lambda}\geqslant\mathrm{ES}_{\Lambda^{\prime}}$ when $\Lambda\geqslant\Lambda^{\prime}$; (b) $\mathrm{ES}_{\Lambda}$ is monotone; (c) $\mathrm{ES}_{\Lambda}\geqslant\mathrm{VaR}_{\Lambda}$; (d) $\mathrm{ES}_{\Lambda}$ is quasi-convex; (e) $\mathrm{ES}_{\Lambda}$ is cash-subadditive; (f) $\mathrm{ES}_{\Lambda}$ is normalized; (g) $\mathrm{ES}_{\Lambda}$ is SSD-consistent; (h) $\mathrm{ES}_{\Lambda}$ is quasi-concave in mixtures; (i) $\mathrm{ES}_{\Lambda}$ is $L^{1}$-continuous when $\Lambda$ takes values in $[0,1)$.
@@
Items (a) and (b) are straightforward because $\mathrm{ES}_{\alpha}(X)$ is monotone (increasing) in both $\alpha$ and $X$, and the supremum of monotone transformations of $\mathrm{ES}_{\alpha}(X)$ is also monotone.
Items (c) and (d) are implied by Theorem 2.
To see item (e), for $c\in\mathbb{R}_{+}$ and $X\in L^{\infty}$, we have
$\mathrm{ES}_{\Lambda}(X+c)$ $=\sup_{x\in\mathbb{R}}\{\mathrm{ES}_{\Lambda(x)}(X+c)\wedge x\}$
$=\sup_{x\in\mathbb{R}}\{(\mathrm{ES}_{\Lambda(x)}(X)+c)\wedge x\}$ [cash additivity of ES]
$=\sup_{x\in\mathbb{R}}\{\mathrm{ES}_{\Lambda(x)}(X)\wedge(x-c)\}+c$
@@
Item (f) follows from (15).
Item (g) follows by applying Lemma 4 of Han et al. (2025), using the fact that $\mathrm{ES}_{\Lambda}$ is cash subadditive, monotone, quasi-convex, and law-invariant. Cash subadditivity is proved in item (e). Law invariance of $\mathrm{ES}_{\Lambda}$ is straightforward from the representation in (12) and the law invariance of ES.
@@
By item (a) of Proposition 2, it is straightforward that $\mathrm{ES}_{\Lambda}$ is bounded above by $\mathrm{ES}_{1}$ and below by $\mathbb{E}$ on $L^{\infty}$. This statement also holds when $\mathrm{ES}_{\Lambda}$ is formulated on larger spaces such as $L^{1}$.
The assumption that $\Lambda$ does not take the value $1$ in item (i) is not dispensable, noting that $\mathrm{ES}_{1}$ is not $L^{1}$-continuous.
We close this section with a result showing that although $\Lambda$-ES is quasi-convex, it is not convex and not concave in mixtures in general, unless it is an ES. This result also highlights the fact that quasi-convexity and convexity are different in strength for cash-subadditive risk measures, although they coincide for monetary risk measures, as shown by Cerreia-Vioglio et al. (2011).
@@
For any decreasing function $\Lambda:\mathbb{R}\to[0,1]$, the following statements are equivalent.
@@
“(iii) $\Rightarrow$ (i)” and “(iii) $\Rightarrow$ (ii)” follow from the facts that ES is convex and ES is concave in mixtures.
“(i) $\Rightarrow$ (iii)”: Suppose that $\mathrm{ES}_{\Lambda}$ is convex and for contradiction that $\Lambda$ is not constant on $\mathbb{R}$. There exist $x>y$ with $\Lambda(x-)<\Lambda((x+y)/2)\leqslant\Lambda(y)$. Take $X,Y\in L^{\infty}$ with $1-\mathbb{P}(X=x)=\mathbb{P}(X=y)=\Lambda((x+y)/2)$ and $Y=y$. It follows that $\mathrm{ES}_{\Lambda}(Y)=y$ and $\mathrm{ES}_{\Lambda}((X+Y)/2)=(x+y)/2$. Because $\Lambda(x-)<\Lambda((x+y)/2)$, we have $\mathrm{ES}_{\Lambda(x-)}(X)<x$. By (14), we have $\mathrm{ES}_{\Lambda}(X)<x$. It follows that $\mathrm{ES}_{\Lambda}(X)/2+\mathrm{ES}_{\Lambda}(Y)/2<\mathrm{ES}_{\Lambda}((X+Y)/2)$, contradicting the convexity of $\mathrm{ES}_{\Lambda}$. Therefore, $\Lambda$ is constant on $\mathbb{R}$.
“(ii) $\Rightarrow$ (iii)”: Suppose that $\Lambda$ is not constant on $\mathbb{R}$. Since $\Lambda$ is bounded, it cannot be concave. Hence, there exist distinct points $x,y,z\in\mathbb{R}$ and $\theta\in(0,1)$ such that $z=\theta x+(1-\theta)y$ and $\Lambda(z)<\theta\Lambda(x)+(1-\theta)\Lambda(y).$ By the continuity of linear functions, there exists $\gamma\in(0,1)$ in any neighborhood of $\theta$ such that $z<\gamma x+(1-\gamma)y$ and $\Lambda(z)<\gamma\Lambda(x)+(1-\gamma)\Lambda(y).$ Write $p=\Lambda(x)$, $q=\Lambda(y)$ and $r=\gamma p+(1-\gamma)q$.
Take independent events $A,B,C\in\mathcal{F}$ such that $\mathbb{P}(A)=1-p$, $\mathbb{P}(B)=1-q$, and $\mathbb{P}(C)=\gamma$. For some constant $K>\max\{-x,-y\}$ (to be determined later), let
$X=x\mathds{1}_{A}-K\mathds{1}_{A^{c}},~~Y=y\mathds{1}_{B}-K\mathds{1}_{B^{c}},\mbox{~~and~~}Z=\mathds{1}_{C}X+\mathds{1}_{C^{c}}Y.$
@@
Note that $\mathbb{P}(Z=-K)=\gamma\mathbb{P}(X=-K)+(1-\gamma)\mathbb{P}(Y=-K)=\gamma p+(1-\gamma)q=r$ and $\Lambda(z)<r$. Therefore,

### v2

@@ 2.3 Expected Shortfall
Hence, we have $\rho(X)\geqslant\mathbb{E}[X]$ in the same sense as the argument in part (i). As $\mathbb{E}$ dominates the essential infimum $\mathrm{VaR}^{+}_{0}$, it implies that $\mathbb{E}$ is the smallest quasi-convex and law-invariant risk measure that dominates $\mathrm{VaR}^{+}_{0}$. ∎
Theorem 1 is stronger than two classical results: Föllmer and Schied (2016, Theorem 4.67), which requires $\rho$ in (9) to be convex, monetary and Fatou-continuous, and Delbaen (2012, Theorem 52), which requires $\rho$ in (9) to be coherent. [^3] We say a risk measure $\rho:\mathcal{X}\to\mathbb{R}$ is Fatou-continuous if it is lower semicontinuous under bounded pointwise convergence: For all bounded $X,X_{1},X_{2},\dots\in\mathcal{X}$ such that $X_{n}\to X$ pointwise as $n\to\infty$, $\rho(X)\leqslant\liminf_{n\to\infty}\rho(X_{n})$. Both of the two results above further assumed that $\rho$ takes finite values and $\alpha\in(0,1)$, but these differences are not essential. We note that (9) fails for $\alpha=0$ because $\mathrm{VaR}_{0}=$-$\infty$ is quasi-convex, and the smallest quasi-convex and law-invariant risk measure dominating $\mathrm{VaR}_{0}$ is itself instead of $\mathrm{ES}_{0}=\mathbb{E}$. Similarly, (10) fails for $\alpha=1$ because $\mathrm{VaR}_{1}^{+}=\infty$ is quasi-convex, and the smallest quasi-convex and law-invariant risk measure dominating $\mathrm{VaR}_{1}^{+}$ is itself instead of $\mathrm{ES}_{1}=\mathrm{VaR}_{1}$.
@@ 3 Lambda ES
Now we turn to our main task of formulating the $\Lambda$-Expected Shortfall ($\Lambda$-ES). By introducing a new class of risk measures, there should be some clear gain. Otherwise, the newly defined class is not useful. The following properties are also satisfied by $\Lambda$-VaR, and they will be considered as basic requirements for $\Lambda$-ES. We believe their desirability is self-evident.
@@
The next four properties are additional requirements for $\Lambda$-ES to be considered a useful alternative to $\Lambda$-VaR, and they highlight the contrasts between ES and VaR.
@@
  • $\Lambda$-ES should be $L^{1}$-continuous. This property models a form of robustness for law-invariant risk measures (Krätschmer et al., 2014).
Some other properties, such as normalization, cash subadditivity, and quasi-concavity in mixtures are also natural from the corresponding properties of ES, although they may be less critical.
With these desirable properties in mind, we are ready to define the $\Lambda$-ES. We first give the formal definition, which is inspired by the $\Lambda$-VaR representtion in Proposition 1, and then show that it satisfies all desirable properties discussed above.
@@
Note the distinction in notation: $\mathrm{ES}_{\Lambda}$ is the mathematical object in (11), whereas $\Lambda$-ES refers to the concept, as we have been speaking of it without formal definition.
Figure 1 illustrates the definition of $\mathrm{ES}_{\Lambda}$ when the function $\Lambda$ is continuous and discontinuous, respectively. Writing $x^{*}=\mathrm{ES}_{\Lambda}(X)$, we can see that $(x^{*},x^{*})$ is the unique intersection point between the graph (linearly interpolated) of the function $x\mapsto\mathrm{ES}_{\Lambda(x)}(X)$ and the graph of the identity. From the right panel of Figure 1, we can also see that whether $x\mapsto\mathrm{ES}_{\Lambda(x)}(X)$ is left- or right-continuous at $x^{*}$ (or neither) does not matter.
@@
By definition, $\mathrm{ES}_{\Lambda}$ is finite on $L^{\infty}$; see also item (a) of Proposition 2 below. The next result shows that $\mathrm{ES}_{\Lambda}$ is the smallest quasi-convex and law-invariant mapping dominating $\mathrm{VaR}_{\Lambda}$. Therefore, it is the unique formulation of the concept of $\Lambda$-ES that generalizes the ES-VaR relation in Theorem 1.
@@
    For a decreasing function $\Lambda:\mathbb{R}\to(0,1]$, the smallest quasi-convex and law-invariant risk measure on $L^{\infty}$ dominating $\mathrm{VaR}_{\Lambda}$ is $\mathrm{ES}_{\Lambda}$, that is, ES_Λ= min{ρ: L^∞→Rρ⩾VaR_Λ and $\rho$ is quasi-convex and law invariant}.
@@
    For a decreasing function $\Lambda:\mathbb{R}\to[0,1)$, the smallest quasi-convex and law-invariant risk measure on $L^{\infty}$ dominating $\mathrm{VaR}^{+}_{\Lambda}$ is $\mathrm{ES}_{\Lambda}$, that is, ES_Λ= min{ρ: L^∞→Rρ⩾VaR^+_Λ and $\rho$ is quasi-convex and law invariant}.
@@
Using (2)–(11), we can see that $\mathrm{ES}_{\Lambda}$ dominates $\mathrm{VaR}_{\Lambda}$ for $\Lambda:\mathbb{R}\to[0,1]$ (resp. $\mathrm{VaR}^{+}_{\Lambda}$ for $\Lambda:\mathbb{R}\to[0,1)$) because $\mathrm{ES}_{\alpha}\geqslant\mathrm{VaR}_{\alpha}$ for all $\alpha\in[0,1]$ (resp. $\mathrm{ES}_{\alpha}\geqslant\mathrm{VaR}^{+}_{\alpha}$ for all $\alpha\in[0,1)$). Moreover, $\mathrm{ES}_{\Lambda}$ is law invariant by definition. Next, we show that $\mathrm{ES}_{\Lambda}$ is quasi-convex. Note that for any given $\alpha\in[0,1]$, $\mathrm{ES}_{\alpha}$ is quasi-convex. Further, an increasing transform of a quasi-convex function is quasi-convex, as well as the supremum of a set of quasi-convex functions. Using these facts,
$\mathrm{ES}_{\Lambda(x)}\mbox{ is quasi-convex for each $x\in\mathbb{R}$}$ $\implies\mathrm{ES}_{\Lambda(x)}\wedge x\mbox{ is quasi-convex for each $x\in\mathbb{R}$}$
@@
Next, for $\Lambda:\mathbb{R}\to(0,1]$, we show that for any $\rho$ that is quasi-convex, law invariant, and satisfying $\rho\geqslant\mathrm{VaR}_{\Lambda}$, it must be $\rho\geqslant\mathrm{ES}_{\Lambda}$. For any $X\in L^{\infty}$, we have:
@@
For $\Lambda:\mathbb{R}\to[0,1)$, for any $\rho$ that is quasi-convex, law invariant, and satisfying $\rho\geqslant\mathrm{VaR}^{+}_{\Lambda}$, we have $\rho(X)\geqslant\mathrm{ES}_{\Lambda}(X)$ for any $X\in L^{\infty}$ with the same argument as above by replacing $\mathrm{VaR}_{\Lambda(x)}$ by $\mathrm{VaR}^{+}_{\Lambda(x)}$ for all $x\in\mathbb{R}$. This completes the proof of statements (i) and (ii). The final statement in (12) follows from (2), by noting that an ES curve $\alpha\mapsto\mathrm{ES}_{\alpha}(X)$ for $X\in L^{\infty}$ can be written as a VaR (resp. $\mathrm{VaR}^{+}$) curve $\alpha\mapsto\mathrm{VaR}_{\alpha}(Y)$ for some $Y\in L^{0}$ on $\alpha\in(0,1]$ (resp. $\alpha\in[0,1)$); see e.g., Lemma 4.5 of Burzoni et al. (2022). ∎
In Theorem 2, the reason to exclude 0 (resp. $1$) in part (i) (resp. part (ii)) from the range of $\Lambda$ is the same as that in Theorem 1, where 0 (resp. $1$) is excluded from the domination of $\mathrm{ES}_{\alpha}$ over $\mathrm{VaR}_{\alpha}$ (resp. $\mathrm{VaR}_{\alpha}^{+}$), explained at the end of Section 2.3.
@@
moreover, if $\Lambda$ is continuous, then (14) becomes an equivalence. These relations will be convenient in some proof arguments.
Remark 1.
As standard in the risk measures literature, the main results are formulated on the space $L^{\infty}$ of essentially bounded random variables. The results in Theorems 1 and 2 hold on the space $L^{1}$ of integrable random variables for $\alpha$ and $\Lambda$ taking values in $(0,1)$, following the same proof arguments, but for $\alpha=0$ and $\alpha=1$, some minor adjustments are needed. It is straightforward to see that all other results hold on $L^{1}$. We discuss the details in Section 6.
@@
It is clear that $\mathrm{ES}_{\Lambda}$ is parameterized only by the function $\Lambda$, and $\mathrm{ES}_{\Lambda}=\mathrm{ES}_{\alpha}$ when $\Lambda\equiv\alpha$ for some $\alpha\in(0,1)$. It satisfies all other desirable properties as a good candidate for $\Lambda$-$\mathrm{ES}$ as discussed in the beginning of the section, which we summarize in the following result.
@@
For any decreasing functions $\Lambda,\Lambda^{\prime}:\mathbb{R}\to[0,1]$, the risk measure $\mathrm{ES}_{\Lambda}$ satisfies the following properties: (a) $\mathrm{ES}_{\Lambda}\geqslant\mathrm{ES}_{\Lambda^{\prime}}$ when $\Lambda\geqslant\Lambda^{\prime}$; (b) $\mathrm{ES}_{\Lambda}$ is monotone; (c) $\mathrm{ES}_{\Lambda}\geqslant\mathrm{VaR}_{\Lambda}$; (d) $\mathrm{ES}_{\Lambda}$ is quasi-convex; (e) $\mathrm{ES}_{\Lambda}$ is normalized; (f) $\mathrm{ES}_{\Lambda}$ is cash subadditive; (g) $\mathrm{ES}_{\Lambda}$ is SSD-consistent; (h) $\mathrm{ES}_{\Lambda}$ is quasi-concave in mixtures; (i) $\mathrm{ES}_{\Lambda}$ is $L^{1}$-continuous when $\Lambda$ takes values in $[0,1)$.
@@
Items (a) and (b) are straightforward because $\mathrm{ES}_{\alpha}(X)$ is monotone (increasing) in both $\alpha$ and $X$, and the supremum of monotone transformations of $\mathrm{ES}_{\alpha}(X)$ is also monotone. Items (c) and (d) are implied by Theorem 2. Item (e) follows from (14).
To see item (f), for $c\in\mathbb{R}_{+}$ and $X\in L^{\infty}$, we have
$\mathrm{ES}_{\Lambda}(X+c)=\sup_{x\in\mathbb{R}}\{\mathrm{ES}_{\Lambda(x)}(X+c)\wedge x\}$ $=\sup_{x\in\mathbb{R}}\{(\mathrm{ES}_{\Lambda(x)}(X)+c)\wedge x\}$ [cash additivity of ES]
@@
Item (g) follows by applying Lemma 4 of Han et al. (2025), using the fact that $\mathrm{ES}_{\Lambda}$ is cash subadditive, monotone, quasi-convex, and law invariant. Cash subadditivity is proved in item (e). Law invariance of $\mathrm{ES}_{\Lambda}$ is clear from the representation in (11) and the law invariance of ES.
@@
By item (a) of Proposition 2, it is straightforward that $\mathrm{ES}_{\Lambda}$ is bounded above by $\mathrm{ES}_{1}$ and below by $\mathbb{E}$ on $L^{\infty}$, which also holds when $\mathrm{ES}_{\Lambda}$ is formulated on larger spaces such as $L^{1}$. The assumption that $\Lambda$ does not take the value $1$ in item (i) is not dispensable, noting that $\mathrm{ES}_{1}$ is not $L^{1}$-continuous.
The next result shows that although $\Lambda$-ES is quasi-concave in mixtures and quasi-convex, it is neither concave in mixtures nor convex in general, unless it is an ES. This result also highlights the fact that quasi-convexity and convexity are different in strength for cash-subadditive risk measures, although they coincide for monetary risk measures, as shown by Cerreia-Vioglio et al. (2011).
@@
For any decreasing function $\Lambda:\mathbb{R}\to[0,1]$, the following are equivalent.
@@
“(iii) $\Rightarrow$ (i)” and “(iii) $\Rightarrow$ (ii)” follow from the facts that ES is convex and ES is concave in mixtures. To prove “(i) $\Rightarrow$ (iii)”, suppose that $\mathrm{ES}_{\Lambda}$ is convex and for contradiction that $\Lambda$ is not constant on $\mathbb{R}$. There exist $x>y$ with $\Lambda(x-)<\Lambda((x+y)/2)\leqslant\Lambda(y)$. Take $X,Y\in L^{\infty}$ with $1-\mathbb{P}(X=x)=\mathbb{P}(X=y)=\Lambda((x+y)/2)$ and $Y=y$. It follows that $\mathrm{ES}_{\Lambda}(Y)=y$ and $\mathrm{ES}_{\Lambda}((X+Y)/2)=(x+y)/2$. Because $\Lambda(x-)<\Lambda((x+y)/2)$, we have $\mathrm{ES}_{\Lambda(x-)}(X)<x$. By (13), we have $\mathrm{ES}_{\Lambda}(X)<x$. It follows that $\mathrm{ES}_{\Lambda}(X)/2+\mathrm{ES}_{\Lambda}(Y)/2<\mathrm{ES}_{\Lambda}((X+Y)/2)$, contradicting the convexity of $\mathrm{ES}_{\Lambda}$. Therefore, $\Lambda$ is constant on $\mathbb{R}$.
“(ii) $\Rightarrow$ (iii)”: Suppose that $\Lambda$ is not constant on $\mathbb{R}$. Since $\Lambda$ is bounded, it cannot be concave. Hence, there exist distinct points $x,y,z\in\mathbb{R}$ and $\theta\in(0,1)$ such that $z=\theta x+(1-\theta)y$ and $\Lambda(z)<\theta\Lambda(x)+(1-\theta)\Lambda(y).$ By the continuity of linear functions, there exists $\gamma\in(0,1)$ in any neighborhood of $\theta$ such that $z<\gamma x+(1-\gamma)y$ and $\Lambda(z)<\gamma\Lambda(x)+(1-\gamma)\Lambda(y).$ Write $p=\Lambda(x)$, $q=\Lambda(y)$ and $r=\gamma p+(1-\gamma)q$. Take independent events $A,B,C\in\mathcal{F}$ such that $\mathbb{P}(A)=1-p$, $\mathbb{P}(B)=1-q$, and $\mathbb{P}(C)=\gamma$. For some constant $K>\max\{-x,-y\}$ (to be determined later), let
$X=x\mathbf{1}_{A}-K\mathbf{1}_{A^{c}},~~Y=y\mathbf{1}_{B}-K\mathbf{1}_{B^{c}},\mbox{~~and~~}Z=\mathbf{1}_{C}X+\mathbf{1}_{C^{c}}Y.$
@@
Because $\mathbb{P}(Z=-K)=\gamma\mathbb{P}(X=-K)+(1-\gamma)\mathbb{P}(Y=-K)=\gamma p+(1-\gamma)q=r$ and $\Lambda(z)<r$,
@@
Remark 3.
Another feature of ES and VaR is that they are tail risk measures in the sense of Liu and Wang (2021). More precisely, for $\alpha\in(0,1)$, an $\alpha$-tail risk measure is a risk measure $\rho$ such that $\rho(X)=\rho(Y)$ when the left quantile functions of $X$ and $Y$ coincide on $(\alpha,1)$. It is straightforward to verify that $\mathrm{ES}_{\Lambda}$ (resp. $\mathrm{VaR}_{\Lambda}$) is an $\alpha$-tail risk measure if and only if $\Lambda\geqslant\alpha$ (resp. $\Lambda>\alpha$) on $\mathbb{R}$.